OpenStudy (luigi0210):

Find the limit:

OpenStudy (luigi0210):

\[\lim_{x \rightarrow 0}~\frac{\sin^{-1}x}{x}\]

geerky42 (geerky42):

I would use L'Hopital's Rule if I were you.

geerky42 (geerky42):

Do you know how to apply L'Hopital's Rule?

OpenStudy (luigi0210):

Yea, just had to remember what \(\frac{d}{dx}sin^{-1}x\) was.. So with L'Hopital I get \(\LARGE \lim_{x \rightarrow 0} \frac{1}{\sqrt{1-x^2}}\)?

geerky42 (geerky42):

Yes, now all you have to do is plug in 0.

OpenStudy (luigi0210):

Alright, thank you!

geerky42 (geerky42):

Glad I helped!